OpenStudy (anonymous):

sin(a+b)=sin a cos b - cos a sin b for all values of a and b. true or false?

myininaya (myininaya):

Why don't you test it for some value a and some value b and see if both sides hold

OpenStudy (anonymous):

False

OpenStudy (anonymous):

sin(a+b)= sinAcosB+cosAsinB

myininaya (myininaya):

See what you get on both sides if you let \[a=\frac{\pi}{2}, b=\frac{\pi}{3}\] You should get a counterexample out of this

OpenStudy (anonymous):

false i think...

myininaya (myininaya):

\[\sin(\frac{5 pi}{6})=? \sin(\frac{\pi}{2})\cos(\frac{\pi}{3})-\cos(\frac{\pi}{2})\sin(\frac{\pi}{2})\] \[\frac{1}{2} =? -1 \cdot \frac{1}{2}-0=\frac{-1}{2} \text{ \not equal}\]

myininaya (myininaya):

Therefore the conclusion is it is not equal for all values a and b

myininaya (myininaya):

So the statement you have above is false